Circular No. 7294

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999em IN NGC 1637 W. D. Li, University of California at Berkeley, on behalf of the Lick

Observatory Supernova Search, report the discovery, on an unfiltered CCD frame taken on Oct. 29.44 UT with the 0.8-m Katzman Automatic Imaging Telescope, of a supernova (mag ∼ 13.5) located at α = 4^{h}41^{m}27^{s}.04, δ = −2^{o}51′45′′.2 (equinox 2000.0), which is 15′′.2 west and 25′′.1 north of the nucleus of NGC 1637. SN 1999em was confirmed at the Beijing Astronomical Observatory on Oct. 29.7. A KAIT image of the same field on Oct. 20.45 showed nothing at the position of SN 1999em (limiting mag ∼ 19.0).

XTE J1859+226 R. I. Hynes, Southampton University; and C. A. Haswell, A. J. Norton,

and S. Chaty, Open University, write: “Following optical and near-infrared monitoring of XTE J1859+226 (IAUC 7279, 7284), Hubble Space Telescope (HST) STIS observations were performed on Oct. 18.10–18.38 UT spanning 112–1026 nm. The ultraviolet spectrum shows broad (12 000 km/s FWZI) and deep Lyα absorption, strong C IV 155-nm emission (EW = 1.1 nm; 4000 km/s FWZI) and weaker emission lines of C III, N V, O III, O IV, O V, Si IV, and He II. Sharp absorption lines of Si II, Si III, Mg II, and other species also appear with possible interstellar origin. Preliminary analysis of the 220-nm interstellar feature suggests E(B − V) = 0.58 ± 0.07. The dereddened optical spectral energy distribution can be represented by a steep blue power law (spectral index +1.2), flattening in the ultraviolet. This resembles energy distributions of other short-period soft-x-ray transients, suggesting that XTE J1859+226 has a binary period < 1 day. A 20-percent decline in 110–170-nm fluxes during the observation coincides with an x-ray decline seen in RXTE/ASM lightcurves. Shorter timescale variations are present, but no persistent coherent modulation was detected. Preliminary analyses of 33 hr of optical time-series data collected between Oct. 13.86 and 20.24 at the Nordic Optical Telescope (J. E. Solheim, R. Ostensen, T. M. C. Abbott), Braeside Observatory (R. Fried, J. McFarland), the Jacobus Kapteyn Telescope (D. J. Rolfe, D. A. Lott), Keele (Z. Ioannou), Mt. Laguna (A. Shafter), St. Andrews (K. O’Brien, K. Horne), and Manastash Ridge Observatory (K. Krisciunas), and also 2.5 hr of infrared time-series data at the U.K. Infrared Telescope (R. J. Ivison) on Oct. 18.19–18.31, suggests a 1-percent modulation with period 22–23 min. Further HST observations are scheduled for Nov. 6.82–6.93 and 19.43–19.55; coordinated groundbased observations are encouraged.”

1999 October 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green